OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). A. y = 2x - 7 B. y=-1/2x-2 C. y = 2x - 2 D. y=-1/2x-7

Mehek (mehek14):

parallel lines = same slope so which answer choices have the same slope as y = 2x + 2 ?

OpenStudy (anonymous):

A and C?

OpenStudy (anonymous):

Here's a previous post with how to do it and why you would do those steps http://openstudy.com/study#/updates/4fd9243be4b091dda540af19

Mehek (mehek14):

correct

Mehek (mehek14):

now we need to find which one of those lines has the points (5,3)

OpenStudy (anonymous):

C?

Mehek (mehek14):

let's start with A \(3=2*5-7\)

Mehek (mehek14):

2*5?

OpenStudy (anonymous):

10

Mehek (mehek14):

10 - 7?

OpenStudy (anonymous):

3

Mehek (mehek14):

\(3=3\) ?

OpenStudy (anonymous):

1?

Mehek (mehek14):

no is 3 = 3 ?

OpenStudy (anonymous):

yes?

OpenStudy (arindameducationusc):

I agree with @Mehek14 Same slope

Mehek (mehek14):

so A is your answer we can check c as well \(y=2x-2\\3=2*5-2\\2*5=10-2=8\\3~is~not~equal~to~8\)

OpenStudy (anonymous):

ohhh thats right lol thnx :)

Mehek (mehek14):

anytime ^_^